# friends.test: rank-based method for feature selection in interaction matrices

### v1

@@ 1 Introduction
Many modern problems involve understanding the interaction between two sets of objects. For instance, recommendation systems link users to movies, pharmacovigilance connects drugs to adverse effects [timilsina2019discovering], and transcriptomics associates genes with biological processes [fertig_cogaps_2010, stein-obrien_enter_2018]. However, not all interactions are equally informative. For instance, in the analysis of protein-protein interaction maps, prioritizing proteins that interact strongly with only a narrow set of biological processes—rather than those with broad, non-specific connectivity—can improve therapeutic specificity and reduce off-target effects [viacava2021centrality]. Genes that are uniformly expressed across all samples do not contribute to the identification of specific biological states. Instead, the analysis relies on tissue-specific markers that provide a clear signal for differentiating unique processes [fertig_cogaps_2010].
Selecting entities for analysis solely by their interaction strength often fails to distinguish true relevance from non-specific background activity, because in many domains meaningful functional relationships are confined to a narrow subset of interactions. For instance, modern genetic studies highlight the necessity of quantifying gene specificity to better identify signals unique to a particular trait [spence2025specificity]. This phenomenon reflects a fundamental challenge in the analysis of bipartite interaction data.
@@ 3 Experimental results
We developed a novel R package friends.test that implements the functionality described above and is available at https://github.com/favorov/friends.test. To validate our friends.test method on the real-world data, we applied it to the previously published transcriptomic dataset (GSE112026) [ando2019chromatin, guo2016characterization]. That dataset contained 47 human papillomavirus-positive head and neck squamous cell carcinoma (HPV+ HNSCC) and 25 normal uvulopharyngoplasty (UPPP) surgical specimens. The method was applied to the RSEM-normalized gene expression matrix to identify the friends. To post-process the data and to interpret the results, we look for group-specific markers among the genes with friends. We say that a gene is a group-specific marker if it has “friends” in at least 25% of the target group samples and has zero “friends” in the opposing group. Based on this criterion, we identified cancer markers (associated exclusively with cancer samples) and normal tissue markers (associated exclusively with normal samples).
@@
To study biological functions of 37 stable markers and the their relationship to cancer, we performed a literature analysis (see Section A), and found that the 35 cancer markers collectively describes an invasive and remodeling tumor phenotype, including the invasion machinery (MMP3, MMP10, MMP13) and extracellular matrix (ECM) (POSTN, SCUBE3, cancer-associated fibroblast (CAF) markers COL11A1, and COL22A1), as well as tumor-specific antigens (CSAG2, PRAME, KRT17, SYCP2). The two normal markers included Chloride Intracellular Channel 3 (CLIC3) and complement C3d receptor 2 (CR2). These results confirm that our friends.test method functions as a high-fidelity biological filter. Section A provides a full list of genes’ functions. Additionally, we have performed GSEA-MSigDB enrichment analysis for the C2 gene set collection [msigdb_genesets] of the cancer marker genes. This analysis showed that those genes were overrepresented mainly in extracellular matrix (ECM) remodeling pathways and in the HNSCC early markers set (see Supplementary File 1).
@@ 3.1 Friends’ set similarity analysis
To illustrate the applicability of the method for assessing functional similarity between cancer gene markers, we utilized the Weighted Jaccard Similarity (also known as Ruzicka Similarity) [cha2007comprehensive]. First, we constructed a global feature space defined by the union of all samples identified across all cancer marker genes. Each gene was then represented as a high-dimensional vector within this space. To characterize these marker genes, we utilized rank-based weighting. Specifically, for a given gene, each sample was assigned a weight based on its rank $r$ (defined as $r^{-1/2}$). Samples without an associated marker were assigned a weight of zero. Next, we quantified the pairwise functional overlap between genes using the Weighted Jaccard index. The resulting similarity matrix served as the input for hierarchical clustering (using the average linkage method [virtanen2020scipy]), see Fig. 2. Section 3.1 discusses the result.
@@
The weighted Jaccard distance matrix (see Fig. 2) contains two pairs of very close genes. The first (and the most close) pair is CXorf59 and CXorf22. Indeed, the up-to-date gene annotation has unified these identifiers under the gene symbol CFAP47 (see [ncbi_gene_170063]). The second pair is LST-3TM12 and SLCO1B3. LST-3TM12 is a legacy identifier for a transcript now classified within the SLCO1B7 genomic region, involved in the same SLCO1B gene family. The two genes are located adjacent on chromosome 12. Moreover, these two genes are sometimes transcribed in the same frame, forming a readthrough transcript protein [uniprot_F5H094]. Notably, the friends.test identifies specific interactions based solely on the internal structure of the input matrix, without relying on external biological databases or pre-existing gene annotations.
@@ 4.1 Related works
Note that the interaction matrix $A$ inherently represents an adjacency matrix of a weighted bipartite graph (network), where two distinct sets of nodes are connected by edges representing their interaction strength. In network analysis, a rich family of methods focuses on identifying globally important nodes (hubs). These approaches prioritize broadly connected entities and capture global importance within a graph structure [ma2017multi, tomasi2011functional]. As interaction datasets grow in complexity, graph anomaly detection has emerged as a critical field. It aims to identify unusual graph instances (nodes, edges, or subgraphs) that deviate significantly from the norm [qiao2025deep, kim2022graph]. A significant area of research involves identifying dense sub-matrices or “blocks” within the interaction matrix, which allows for detecting structural changes at the matrix or submatrix level. Specifically, biclustering methods focus on discovering sub-matrices that satisfy specific homogeneity and statistical significance criteria [castanho2024biclustering, jose2022biclustering, pontes2015biclustering].
Our approach operates at the individual profile level (i.e., row-wise), rather than attempting to partition the entire interaction matrix at once. However, even at the row level, the choice of method remains critical. While clustering techniques like $k$-means are widely used for partitioning data, their application in our case presents some limitations. Specifically, $k$-means assumes relatively balanced cluster sizes and can be sensitive to outliers [zhou2024adaptive, klochkov2021robust], which may be the case when the number of “friendly” interactions and background interactions differ significantly.
@@
Widely used specificity indices, such as the Tau index or the Gini coefficient, quantify how unevenly values are distributed within an interaction profile, thereby providing a single-number measure of overall specificity. However, these metrics do not identify which particular interactions constitute the specific signal [kryuchkova2017benchmark].
@@ 37 CR2 Interface between innate and adaptive immune systems
  • COL11A1 is widely validated as a specific marker for CAFs in the head and neck tumor microenvironment [nallanthighal2021collagen, garcia2013overexpression].
  • MMP3, MMP10, and MMP13 are critical for degrading the basement membrane, facilitating tumor invasion. Specifically, MMP10 and MMP13 are known to correlate with metastasis and poor survival in HNSCC [deraz2011mmp, luukkaa2006association, vincent2014overexpression, iizuka2014matrix].
  • [wang2023machine] identified COL7A1 as a top-ranking diagnostic predictor specifically for squamous cell carcinomas, including Head and Neck Squamous Cell Carcinoma (HNSC) and Lung Squamous Cell Carcinoma (LUSC).
  • POSTN (Periostin) functions as a hub gene for cell adhesion and migration, bridging cancer cells with the structural matrix [teng2025periostin, zhao2022hub, xian2025periostin].
  • Secretory SCUBE3 supports oncogenic activity through interactions with key oncogenic cell surface receptor proteins [singh2025antibody].
  • PRAME is highly specific to HNSCC and melanoma and is associated [szczepanski2013prame] with retinoid resistance [ramchatesingh2025targeting]
  • KRT17 is identified as a critical mediator of drug resistance and immune evasion in HNSCC [hou2025krt17].
  • SYCP2 is expressed in HPV associated Cancers [almubarak2021investigating].

### v2

@@ 1 Introduction
Many modern problems involve understanding the interaction between two sets of objects. For instance, recommendation systems link users to movies, pharmacovigilance connects drugs to adverse effects [28], and transcriptomics associates genes with biological processes [7, 25]. However, not all interactions are equally informative. For instance, in the analysis of protein-protein interaction maps, prioritizing proteins that interact strongly with only a narrow set of biological processes—rather than those with broad, non-specific connectivity—can improve therapeutic specificity and reduce off-target effects [31]. Genes that are uniformly expressed across all samples do not contribute to the identification of specific biological states. Instead, the analysis relies on tissue-specific markers that provide a clear signal for differentiating unique processes [7].
Selecting entities for analysis solely by their interaction strength often fails to distinguish true relevance from non-specific background activity, because in many domains meaningful functional relationships are confined to a narrow subset of interactions. For instance, modern genetic studies highlight the necessity of quantifying gene specificity to better identify signals unique to a particular trait [24]. This phenomenon reflects a fundamental challenge in the analysis of bipartite interaction data.
@@ 3 Experimental results
We developed a novel R package friends.test that implements the functionality described above and is available at https://github.com/favorov/friends.test. To validate our friends.test method on the real-world data, we applied it to the previously published transcriptomic dataset (GSE112026) [2, 9]. That dataset contained 47 human papillomavirus-positive head and neck squamous cell carcinoma (HPV+ HNSCC) and 25 normal uvulopharyngoplasty (UPPP) surgical specimens. The method was applied to the RSEM-normalized gene expression matrix to identify the friends. To post-process the data and to interpret the results, we look for group-specific markers among the genes with friends. We say that a gene is a group-specific marker if it has “friends” in at least 25% of the target group samples and has zero “friends” in the opposing group. Based on this criterion, we identified cancer markers (associated exclusively with cancer samples) and normal tissue markers (associated exclusively with normal samples).
@@
To study biological functions of 37 stable markers and the their relationship to cancer, we performed a literature analysis (see Section A), and found that the 35 cancer markers collectively describes an invasive and remodeling tumor phenotype, including the invasion machinery (MMP3, MMP10, MMP13) and extracellular matrix (ECM) (POSTN, SCUBE3, cancer-associated fibroblast (CAF) markers COL11A1, and COL22A1), as well as tumor-specific antigens (CSAG2, PRAME, KRT17, SYCP2). The two normal markers included Chloride Intracellular Channel 3 (CLIC3) and complement C3d receptor 2 (CR2). These results confirm that our friends.test method functions as a high-fidelity biological filter. Section A provides a full list of genes’ functions. Additionally, we have performed GSEA-MSigDB enrichment analysis for the C2 gene set collection [3] of the cancer marker genes. This analysis showed that those genes were overrepresented mainly in extracellular matrix (ECM) remodeling pathways and in the HNSCC early markers set (see Supplementary File 1).
@@ 3.1 Friends’ set similarity analysis
To illustrate the applicability of the method for assessing functional similarity between cancer gene markers, we utilized the Weighted Jaccard Similarity (also known as Ruzicka Similarity) [5]. First, we constructed a global feature space defined by the union of all samples identified across all cancer marker genes. Each gene was then represented as a high-dimensional vector within this space. To characterize these marker genes, we utilized rank-based weighting. Specifically, for a given gene, each sample was assigned a weight based on its rank $r$ (defined as $r^{-1/2}$). Samples without an associated marker were assigned a weight of zero. Next, we quantified the pairwise functional overlap between genes using the Weighted Jaccard index. The resulting similarity matrix served as the input for hierarchical clustering (using the average linkage method [33]), see Fig. 2. Section 3.1 discusses the result.
@@
The weighted Jaccard distance matrix (see Fig. 2) contains two pairs of very close genes. The first (and the most close) pair is CXorf59 and CXorf22. Indeed, the up-to-date gene annotation has unified these identifiers under the gene symbol CFAP47 (see [19]). The second pair is LST-3TM12 and SLCO1B3. LST-3TM12 is a legacy identifier for a transcript now classified within the SLCO1B7 genomic region, involved in the same SLCO1B gene family. The two genes are located adjacent on chromosome 12. Moreover, these two genes are sometimes transcribed in the same frame, forming a readthrough transcript protein [30]. Notably, the friends.test identifies specific interactions based solely on the internal structure of the input matrix, without relying on external biological databases or pre-existing gene annotations.
@@ 4.1 Related works
Note that the interaction matrix $A$ inherently represents an adjacency matrix of a weighted bipartite graph (network), where two distinct sets of nodes are connected by edges representing their interaction strength. In network analysis, a rich family of methods focuses on identifying globally important nodes (hubs). These approaches prioritize broadly connected entities and capture global importance within a graph structure [17, 29]. As interaction datasets grow in complexity, graph anomaly detection has emerged as a critical field. It aims to identify unusual graph instances (nodes, edges, or subgraphs) that deviate significantly from the norm [21, 13]. A significant area of research involves identifying dense sub-matrices or “blocks” within the interaction matrix, which allows for detecting structural changes at the matrix or submatrix level. Specifically, biclustering methods focus on discovering sub-matrices that satisfy specific homogeneity and statistical significance criteria [4, 12, 20].
Our approach operates at the individual profile level (i.e., row-wise), rather than attempting to partition the entire interaction matrix at once. However, even at the row level, the choice of method remains critical. While clustering techniques like $k$-means are widely used for partitioning data, their application in our case presents some limitations. Specifically, $k$-means assumes relatively balanced cluster sizes and can be sensitive to outliers [37, 14], which may be the case when the number of “friendly” interactions and background interactions differ significantly.
@@
Widely used specificity indices, such as the Tau index or the Gini coefficient, quantify how unevenly values are distributed within an interaction profile, thereby providing a single-number measure of overall specificity. However, these metrics do not identify which particular interactions constitute the specific signal [15].
@@ 37 CR2 Interface between innate and adaptive immune systems
  • COL11A1 is widely validated as a specific marker for CAFs in the head and neck tumor microenvironment [18, 8].
  • MMP3, MMP10, and MMP13 are critical for degrading the basement membrane, facilitating tumor invasion. Specifically, MMP10 and MMP13 are known to correlate with metastasis and poor survival in HNSCC [6, 16, 32, 11].
  • [34] identified COL7A1 as a top-ranking diagnostic predictor specifically for squamous cell carcinomas, including Head and Neck Squamous Cell Carcinoma (HNSC) and Lung Squamous Cell Carcinoma (LUSC).
  • POSTN (Periostin) functions as a hub gene for cell adhesion and migration, bridging cancer cells with the structural matrix [27, 36, 35].
  • Secretory SCUBE3 supports oncogenic activity through interactions with key oncogenic cell surface receptor proteins [23].
  • PRAME is highly specific to HNSCC and melanoma and is associated [26] with retinoid resistance [22]
  • KRT17 is identified as a critical mediator of drug resistance and immune evasion in HNSCC [10].
  • SYCP2 is expressed in HPV associated Cancers [1].